(i) Prove by mathemarical induction on n that

3n ≥ 2n2 + 1 for all n ∈ N

(ii) Given the function g : R → R defined by

􏰂 x−1

g (x) = 2x+4 1

2

if x̸=−2 if x=−2

Find whether or not f is injective and surjective.

Find the inverse of f, if it exists.


1
Expert's answer
2021-04-28T08:20:48-0400

(i) "3n\ge 2n^2+1"


at "n=1, 3\ge 2+1=3"


P(1) is true


"\text{At } n=k, P(k): 3k\ge 2k^2+1~~~~~-(1)"


at n=k+1,

"p(k+1)" :"3(k+1)\ge 2(k+1)^2+1"


"3k+3\ge 2k^2+2+4k+1"

"3k\ge 2k^2+4k"


As from eqn.(1)-"3k\ge 2k^2+1 \text{ and }3k\ge 2k^2+4k"


"\Rightarrow" The given statement "p(n):3k\ge 2n^2+1" is true "\forall n\in N" .


(ii)

"g(x)=2x+4"


for Every value of x, There are different value g(x) So g(x) is injective.


Since the image im(X) of f equals the codomain function g(x), So g(x) is surjective.

Hence g(x) is injective and surjective.


Inverse:

Let "g(x)=y\Rightarrow 2x+4=y\Rightarrow x=\dfrac{y-4}{2}"


Hence The inverse is- "g^{-1}(x)=\dfrac{x-4}{2}"